OpenStudy (anonymous):

Boron has two isotopes, one with a mass of 10 and the other with a mass of 11. Given that boron has an atomic mass of 10.8, what is the fractional abundance of 10B?

OpenStudy (***[isuru]***):

Hi, This is an easy one The given value for Boron (10.8) is an average valu calculate using the mass of 11B and 10B according to their abundance 10.8 =( percentage of abundance of 10B * 10) +( percentage of abundance of 11B *11) if 10B's abundance is x % , the abundance of 11B is (100 - x) % using that theorem 10.8 = [(x/100) * 10 ]+[ (100-x)/100 * 11] 1080 = 10x + 1100 - 11x x = 20 that's it !!! 10B is abundance is 20% which is 1/5 Hope this will help u out